# LLMs on a Budget? Say HOLA

### v1

@@ Abstract
Running Large Language Models (LLMs) on edge devices is constrained by high compute and memory demands—posing a barrier for real-time applications in sectors like healthcare, education, and embedded systems. Current solutions such as quantization, pruning, and retrieval-augmented generation (RAG) offer only partial optimizations and often compromise on speed or accuracy.
We introduce HOLA, an end-to-end optimization framework for efficient LLM deployment. Internally, it leverages Hierarchical Speculative Decoding (HSD) for faster inference without quality loss. Externally, AdaComp-RAG adjusts retrieval complexity based on context needs. Together with Lo-Bi, which blends structured pruning (LoRA) and quantization, HOLA delivers significant gains: +17.6% EMA on GSM8K, +10.5% MCA on ARC, and reduced latency and memory on edge devices like Jetson Nano—proving both scalable and production-ready.
@@ 1 Introduction
Large Language Models (LLMs) have transformed NLP applications—from question answering Ehtesham et al. (2025) and code generation Zhong and Wang (2024) to summarization Ghosh et al. (2024) and conversational agents Zenimoto (2024). However, their deployment on edge and low-resource environments remains constrained by high compute and memory demands Dutta et al. (2025), limiting real-world impact in domains like healthcare Lysandrou et al. (2023), education Bayarri-Planas et al. (2025), and personalized assistants Tack et al. (2024).
While techniques such as quantization Tack et al. (2024), pruning Chimoto et al. (2024), and retrieval-augmented generation (RAG) Parekh et al. (2024) offer partial relief, they often sacrifice accuracy, speed, or generality. Moreover, most approaches optimize isolated components—either internal inference Liu et al. (2023) or external retrieval Zheng et al. (2023)—without addressing the full pipeline.
We introduce HOLA (Hierarchical Optimized Language Augmentation), a unified framework targeting both internal and external efficiency. It integrates (1) Hierarchical Speculative Decoding (HSD) for faster, accurate inference; (2) AdaComp-RAG, an adaptive retrieval system based on context relevance; and (3) Lo-Bi Optimization, combining LoRA-based structured pruning with mixed-precision quantization.
2 Related Work
Deploying LLMs in constrained environments has driven advances in decoding acceleration, retrieval optimization, and model compression. Techniques like Speculative Decoding Sun et al. (2024) and Medusa Cai et al. (2024) use draft-and-verify schemes to speed up inference, though they often require extra verifier networks and struggle with long sequences. Adaptive retrieval methods Labruna et al. (2024) selectively route queries to improve efficiency but introduce added system complexity and retraining needs.
Model compression remains central to edge deployment. Quantization approaches like GPTQ Bumgardner et al. (2025) and Bi-LLM Weng et al. (2025) reduce precision for memory savings, but aggressive quantization can impact accuracy. Structured pruning using LoRA Hu et al. and LoRA-Pruner Zhang et al. (2024b) compress model weights with minimal retraining. While a few efforts combine pruning and quantization Ali et al. (2025), most lack holistic, hardware-aware integration across the inference stack.
@@ 3 Methodology
Our proposed framework, HOLA, is composed of three synergistic modules—HSD accelerates autoregressive generation via entropy-aware verification, AdaComp-RAG adaptively modulates retrieval granularity, and Lo-Bi Optimization ensures compact and resource-aware model deployment. The broader process of the HOLA is illustrated in Algorithm 1.
@@ 3.1 HSD Module
This reduces the number of verifier invocations from $O(T)$ to $O(k)$, where $k\ll T$, enabling up to 2–3$\times$ decoding speedup without significant degradation in output quality. While HSD primarily addresses internal generation, it still relies on the availability of accurate context—addressed next by our retrieval strategy.
@@ 4.2 Evaluation Metrics
4.3 Hyperparameters
The HOLA framework employs a finely tuned configuration across all modules to balance performance and efficiency. In the HSD module, we set the entropy threshold to $\tau=1.5$, minimizing verifier calls without sacrificing quality. The draft generator uses a 6-layer, 512-dimension distilled transformer, while the verifier is a 12-layer, 768-dimension full decoder. AdaComp-RAG uses a retrieval threshold $\delta=0.85$, informed by gradient norms on a validation set. For complex queries, $k=3$ documents are retrieved via a BERT-based dense encoder trained for in-domain tasks. Compositional attention runs with a shared hidden size of 768. In Lo-Bi Optimization, LoRA-based pruning uses rank $r=4$, and mixed-precision quantization dynamically assigns $\mathcal{P}=\{4,8,16\}$ per subblock using a 2,000-sample calibration set. Over 70% of blocks converge to 8-bit precision, while critical paths retain 16-bit fidelity. Training across modules uses the AdamW optimizer (learning rate $2\times 10^{-4}$, batch size 32) with linear warm-up (10%) and early stopping based on EMA (GSM8K) and MCA (ARC). Unless specified, all experiments run on NVIDIA A100 GPUs (80 GB). Code will be released post-acceptance.
5 Experimental Analysis
5.1 Comparison with Baselines
@@ 5.3 Ablation Study
Short ($<$30 tokens HOLA 83.4 102 694
@@
Long ($>$60 tokens) HOLA 89.0 134 720
@@ 6 Additional Analysis
We segment the test set into three bins based on token length: Short ($<$30), Medium (30–60), and Long ($>$60), each with $\sim$300 examples. Table 4 highlights HOLA consistent performance across varying input lengths on GSM8K. Accuracy improves with input length, peaking at 89.0% EMA for long inputs. Despite slight increases in latency and memory, HOLA maintains an efficient tradeoff, demonstrating strong scalability and robustness for complex, multi-hop reasoning tasks.
@@
We benchmark HOLA across edge devices (Jetson Nano, Raspberry Pi 4), general-purpose CPUs (Intel i7-10700K), and high-performance GPUs (NVIDIA A100) using the GSM8K dataset to evaluate deployment feasibility under diverse hardware constraints. As shown in Table 6, HOLA retains high accuracy across platforms—Jetson Nano (87.6%), Raspberry Pi 4 (85.1%), and A100 (89.2%)—demonstrating effective quantization and low-bit inference strategies. Latency scales with compute capability: A100 achieves the fastest inference at 68 ms/sample, outperforming Jetson Nano ($841$ ms) and Raspberry Pi 4 ($1092$ ms). Memory usage remains within edge-operable bounds: Jetson Nano and Raspberry Pi peak at $\sim$1.2 GB, while A100 benefits from efficient batching ($\sim$942 MB). Power consumption is minimal on edge devices (7.5 W and 6.1 W), supporting low-power deployments like IoT and robotics. However, their sub-1.5 samples/sec throughput limits real-time scalability. A100 delivers $14+$ samples/sec at 97.2 W, suiting high-volume inference and server-based workloads. These results validate HOLA’s adaptability to both edge and cloud contexts, enabling flexible deployment strategies based on latency, accuracy, and power trade-offs.
We focus our additional analysis on the GSM8K dataset rather than ARC because GSM8K presents more diverse question structures and longer multi-hop arithmetic reasoning challenges, offering a richer evaluation of HOLA’s ability to handle increasing input complexity and retrieval overhead. Its numerical nature also makes it more sensitive to precision, latency, and memory trade-offs—critical for assessing real-world deployment viability across constrained and high-performance platforms.
@@ 7 Conclusion and Future Work
While HOLA delivers strong performance-efficiency trade-offs, several limitations remain. The retrieval pipeline depends on static indexing, limiting adaptability to dynamic knowledge or temporal queries. Aggressive compression, while beneficial for edge deployment, can introduce non-trivial accuracy loss on ultra-low-memory devices. Current evaluation is constrained to GSM8K, and generalizability to more complex, multi-turn, or multi-modal tasks remains unverified. Robustness under adversarial or noisy retrievals is also underexplored. Additionally, real-world deployment factors—such as thermal throttling, energy spikes, or long-term device wear—are not modeled. Future work will explore adaptive retraining, retrieval refresh strategies, and system-level profiling for deployment-aware optimization.

### v2

@@ Abstract
Running Large Language Models (LLMs) on edge devices is constrained by high compute and memory demands—posing a barrier for real-time applications in industries like healthcare, education, and embedded systems. Current solutions such as quantization, pruning, and Retrieval-Augmented Generation (RAG) offer only partial optimizations and often compromise on speed or accuracy. We introduce HOLA, an end-to-end optimization framework for efficient LLM deployment. Internally, it leverages Hierarchical Speculative Decoding (HSD) for faster inference without quality loss. Externally, AdaComp-RAG adjusts retrieval complexity based on context needs. Together with Lo-Bi, which blends structured pruning (LoRA) and quantization, HOLA delivers significant gains: +17.6% EMA on GSM8K, +10.5% MCA on ARC, and reduced latency and memory on edge devices like Jetson Nano—proving both scalable and production-ready. Our code is available at: https://github.com/zohaibhasan066/HOLA_Codebase
@@ 1 Introduction
Large Language Models (LLMs) have transformed NLP applications—from question answering Ehtesham et al. (2025) and code generation Zhong and Wang (2024) to summarization Ghosh et al. (2024) and conversational agents Zenimoto (2024). However, their deployment on edge and low-resource environments remains constrained by high compute and memory demands Dutta et al. (2025), limiting real-world impact in domains like healthcare Lysandrou et al. (2023), education Bayarri-Planas et al. (2025), and personalized assistants Tack et al. (2024). While techniques such as quantization Tack et al. (2024), pruning Chimoto et al. (2024), and Retrieval-Augmented Generation (RAG) Parekh et al. (2024) offer partial relief, they often sacrifice accuracy, speed, or generality. Moreover, most approaches optimize isolated components—either internal inference Liu et al. (2023) or external retrieval Zheng et al. (2023)—without addressing the full pipeline. We introduce HOLA (Hierarchical Optimized Language Augmentation), a unified framework targeting both internal and external efficiency. It integrates (1) Hierarchical Speculative Decoding (HSD) for faster, accurate inference; (2) AdaComp-RAG, an adaptive retrieval system based on context relevance; and (3) Lo-Bi Optimization, combining LoRA-based structured pruning with mixed-precision quantization.
2 Related Works
Deploying LLMs in constrained environments has driven advances in decoding acceleration, retrieval optimization, and model compression. Techniques like Speculative Decoding Sun et al. (2024) and Medusa Cai et al. (2024) use draft-and-verify schemes to speed up inference, though they often require extra verifier networks and struggle with long sequences. Adaptive retrieval methods Labruna et al. (2024) selectively route queries to improve efficiency but introduce added system complexity and retraining needs. Model compression remains central to edge deployment. Quantization approaches like GPTQ Bumgardner et al. (2025) and Bi-LLM Weng et al. (2025) reduce precision for memory savings, but aggressive quantization can impact accuracy. Structured pruning using LoRA Hu et al. and LoRA-Pruner Zhang et al. (2024b) compress model weights with minimal retraining. While a few efforts combine pruning and quantization Ali et al. (2025), most lack holistic, hardware-aware integration across the inference stack.
Figure 1: HOLA architecture.
@@ 3 Methodology
Our proposed framework, HOLA, is composed of three synergistic modules—HSD accelerates autoregressive generation via entropy-aware verification, AdaComp-RAG adaptively modulates retrieval granularity, and Lo-Bi Optimization ensures compact and resource-aware model deployment. The broader process of the HOLA is illustrated in Algorithm 1. On the other hand, Figure 1 presents the detailed architecture.
@@ 3.1 HSD Module
[…]
@@ 3.3 Lo-Bi Optimization Module
4 Experimental Setup
4.1 Datasets
@@
4.2 Evaluation Metrics
4.3 Hyperparameters
The HOLA framework employs a finely tuned configuration across all modules to balance performance and efficiency. In the HSD module, we set the entropy threshold to $\tau=1.5$, minimizing verifier calls without sacrificing quality. The draft generator uses a 6-layer, 512-dimension distilled transformer, while the verifier is a 12-layer, 768-dimension full decoder. AdaComp-RAG uses a retrieval threshold $\delta=0.85$, informed by gradient norms on a validation set. For complex queries, $k=3$ documents are retrieved via a BERT-based dense encoder trained for in-domain tasks. Compositional attention runs with a shared hidden size of 768. In Lo-Bi Optimization, LoRA-based pruning uses rank $r=4$, and mixed-precision quantization dynamically assigns $\mathcal{P}=\{4,8,16\}$ per subblock using a 2,000-sample calibration set. Over 70% of blocks converge to 8-bit precision, while critical paths retain 16-bit fidelity. Training across modules uses the AdamW optimizer (learning rate $2\times 10^{-4}$, batch size 32) with linear warm-up (10%) and early stopping based on EMA (GSM8K) and MCA (ARC). Unless specified, all experiments run on NVIDIA A100 GPUs (80 GB).
5 Experimental Analysis
5.1 Comparison with Baselines
@@ 5.2 Cross-Domain Generalization Analysis
Short (<<30 tokens HOLA 83.4 102 694
@@
Long (>>60 tokens) HOLA 89.0 134 720
@@ 6 Additional Analysis
We segment the test set into three bins based on token length: Short (<<30), Medium (30–60), and Long (>>60), each with $\sim$300 examples. Table 4 highlights HOLA consistent performance across varying input lengths on GSM8K. Accuracy improves with input length, peaking at 89.0% EMA for long inputs. Despite slight increases in latency and memory, HOLA maintains an efficient tradeoff, demonstrating strong scalability and robustness for complex, multi-hop reasoning tasks.
@@
We evaluate HOLA on GSM8K across Jetson Nano, Raspberry Pi 4, Intel i7-10700K, and NVIDIA A100 to assess deployment across diverse hardware. As shown in Table 6, HOLA maintains high accuracy—87.6% (Jetson Nano), 85.1% (Raspberry Pi 4), and 89.2% (A100)—demonstrating robust low-bit inference. Latency scales with compute: A100 is fastest (68 ms/sample), followed by Jetson Nano (841 ms) and Raspberry Pi 4 (1092 ms). Memory usage remains edge-compatible ($\sim$1.2 GB), and power stays low (7.5 W, 6.1 W) on edge devices. A100 delivers $14+$ samples/sec (97.2 W), suiting high-throughput workloads. HOLA proves deployable across edge and cloud, balancing speed, accuracy, and power. We focus on GSM8K due to its diverse, multi-hop numerical reasoning demands, offering a better testbed for latency and precision trade-offs than ARC.
(a) HOLA effect on model ranking
(b) Transfer: ARC to GSM8K efficiency
(c) Transfer: GSM8K to ARC efficiency
(d) Ranking shifts due to HOLA
(e) GSM8K latent space
(f) ARC latent space
Figure 2: Detailed analysis of the impact of HOLA optimization and domain transfer on large language models (LLMs) across tasks and hardware settings. (a) shows how HOLA influences model ranking for the GSM8K and ARC datasets. (b) and (c) illustrate the efficiency changes when transferring models between ARC and GSM8K tasks on various hardware. (d) compares ranking shifts caused by HOLA optimization. (e) and (f) provide t-SNE visualizations of the latent space to highlight task domain effects for GSM8K and ARC, respectively.
Qualitative Analysis
Figure 2 provides a comprehensive qualitative analysis of the impact of HOLA optimization and domain transfer on LLMs across two benchmark tasks: ARC and GSM8K. Subfigure (a) illustrates how HOLA affects model rankings, showing that the optimization has a varied impact across different models and tasks—some models improve significantly, while others decline or remain unaffected. Subfigures (b) and (c) show the efficiency of cross-domain transfer from ARC to GSM8K and vice versa, respectively. The heatmaps reveal that transfer efficiency is highly dependent on model architecture and hardware configurations, with some models exhibiting strong generalization across tasks and others struggling. Subfigure (d) further highlights the shifts in model rankings caused by HOLA, indicating that optimization can reconfigure relative model performance on these benchmarks. Subfigures (e) and (f) use t-SNE to visualize the latent task spaces of GSM8K and ARC, respectively. The distinct clustering in each plot suggests that the two tasks occupy separate regions in the latent space, which helps explain the varying success of cross-task transfer. Together, these insights emphasize that both optimization and evaluation strategies must consider task specificity and domain effects to ensure robust performance across reasoning benchmarks.
@@ 7 Conclusion and Future Work
While HOLA delivers strong performance-efficiency trade-offs, several limitations remain. The retrieval pipeline depends on static indexing, limiting adaptability to dynamic knowledge or temporal queries. Aggressive compression, while beneficial for edge deployment, can introduce non-trivial accuracy loss on ultra-low-memory devices. Current evaluation is constrained to GSM8K, and generalizability to more complex, multi-turn, or multi-modal tasks remains unverified. Robustness under adversarial or noisy retrievals is also underexplored. Additionally, real-world deployment factors—such as thermal throttling, energy spikes, or long-term device wear—are not modeled.